# DurFlex-EVC: Duration-Flexible Emotional Voice Conversion with Parallel Generation

### v3

@@ V-A Comparison of evaluation results for baseline models
V-B Experiments for Analyzing Model Architectures
V-B1 Effectiveness of Style Autoencoder
We conducted an experiment to verify the effectiveness of the style autoencoder (SAE). w/o SAE is a model that stacks a standard feed forward transformer block with a full layer of SAE, without purtubation and conditioning for emotion style. In Table V, w/o SAE shows the results of the ablation study on it. The overall performance degradation observed in models without SAE indicates the importance of disentangling content and style from input features. Fig. 6 shows the change in cross entropy for styles as input features pass through SAE. To obtain cross entropy, we trained a classifier for emotion style on features extracted from all layers of SAE. The black line is the result for the model with SAE and the orange line is the result for the model without SAE. The w/o SAE model shows high cross entropy across the layers, while the model with SAE shows increasing cross entropy as it passes through the de-stylize transformer layers and decreasing cross entropy as it passes through the stylize transformer layers. This means that the style is disentangled and conditioned from the feature by SAE.
We also experimented with adversarial training strategies to remove source styles from input features. The results for this are shown in Table V as w/ adv, and show a performance degradation on all metrics except EECS. This indicates that the proposed method is more stable for model learning than traditional adversarial learning.
@@ V-B1 Effectiveness of Style Autoencoder
V-B2 Effectiveness of Unit Aligner
The unit aligner (UA) was introduced to model stylized contexts. Table V w/o UA shows the results of the ablation experiment for this. The results showed that w/o UA performed better in terms of voice quality and pronunciation, but not in terms of emotion conversion. This indicates that there remains a lot of information about the source style in the feature and that the UA functions as a bottleneck and has a significant impact on style control. We measured the Jensen-Shannon divergence (JSD) to validate that the predicted units in UA reflect the target emotion. Table VI shows the JSD results of the units for each emotion in Textless-EVC and DurFlex-EVC. DurFlex-EVC achieved a better JSD than Textless-EVC for all emotions, indicating that it provides an appropriate representation of the context.
V-B3 Effectiveness of Hierarchical Stylize Encoder
The output of the unit aligner is a representation in units, which is expanded to frame-level by a duration predictor and length regulator. For frame-level stylization in this representation, we introduce an heirarchical stylize encoder (HSE), and the results of the ablation study are shown in Table V w/o HSE. The evaluation results showed an overall performance degradation with w/o HSE. This is inferred to be due to HSE reducing the burden on the Mel-spectrogram generator, resulting in improved generation quality.
(a) Angry
(b) Happy
(c) Sad
(d) Surprise
@@ V-B3 Effectiveness of Hierarchical Stylize Encoder
V-B4 Comparison for Duration Predictors
We introduced a stochastic duration predictor (SDP) for duration modeling to represent the diversity of emotions. To evaluate, we compared our experiments with widely used deterministic duration predictors (DDP). The DDP follows the structure described in FastSpeech [43], which includes two convolutional layers, ReLU activation, layer normalization, and dropout. Tables V, w/ DDP show the evaluation results of the model using DDP. The evaluation results show that the model using SDP is better than the model using DDP. We compared the JSD of unit duration for each emotion for each duration predictor. Table VII shows the JSD results for unit duration for models with stochastic duration predictors, w/ SDP, and deterministic duration predictors, w/ DDP.
(a) DurFlex-EVC w/ FFT
(b) DurFlex-EVC w/ diffusion
@@ V-B5 Comparison for Mel-spectrogram Generator Structures
To verify the expressiveness of each emotion, we computed the JSD over pitch for each emotion. In Table VIII, w/ FFT is the model using an FFT decoder and w/ diffusion is the model using a diffusion-based structure. w/ diffusion shows better JSD on neutral, angry, sad, while w/ FFT shows better JSD in happy and surprise. Although w/ FFT outscores w/ diffusion in Happy and Surprise, it does not mean that FFT is more expressive. Fig. 8 shows the pitch of the speech generated by multiple iterations of the same sentence. Models using FFT produce consistent pitch tracks that remain stable over repeated generations, while models using diffusion produce more dynamic results. It can be interpreted that the diffusion-based structure models a higher expressiveness, resulting in a higher JSD than the FFT-based structure, which models the average over dynamic emotions (happy, sad).
@@
V-B6 Comparison for Input Features
The proposed model takes HuBERT features as input and outputs a Mel-spectrogram. The reason for outputting Mel-spectrogram is for compatibility with pre-trained vocoders. To explain the reasoning behind our choice of HuBERT features, we present a comparison of the input features. We compared Mel-spectrogram and linear spectrogram as features using conventional signal processing, and wav2vec 2.0 [27], wavLM [32], and HuBERT [29] as SSL features that have recently been used as linguistic representations [35, 33]. Table IX shows the results of the experiments by input feature. The results show that models using Mel-spectrogram and linear-spectrogram have lower UTMOS than SSL feature models. Models using wavLM performed better overall than those using wav2vec 2.0. The model using HuBERT outperformed the others in all metrics except ECA and EECS. In particular, the model using HuBERT shows a significant improvement in pronunciation, which is interpreted as advantageous for linguistic learning over other features because the target units used for training are obtained from the clustering of HuBERT features. To compare the linguistic modeling ability of each input feature, we calculated the BLEU score [72] and the unit error rate (UER) of the prediction unit. Table X shows the BLEU score and UER of the predicted units for each feature. We found that the prediction of units was correlated with the pronunciation accuracy of the synthesized speech and that this was due to the fact that HuBERT predicted units more accurately than the other features.
We also experimented with a model that takes unit input and generates speech. In Table V, w/ unit2mel is a model that receives unit input and generates Mel-spectrogram. The w/ unit2wav is a model that predicts units and generates waveforms. The w/ unit2mel shows a poorer overall performance, except for ECA and EECS. The w/ unit2wav shows significantly worse metrics across the board, except for CER and WER. We interpret this to mean that the speech unit has enough information about pronunciation but not enough other speech information to generate a waveform. To overcome this, Textless-EVC uses additional information such as pitch and timbre to generate the waveform.
Figure 9: Style embedding modeling for unseen speaker setting.
@@ V-B6 Comparison for Input Features
V-C Unseen Speaker Emotion Conversion
Table XI shows the evaluation results for seen and unseen speakers for the modified model. The modified model scored better UTMOS, PER, CER, and WER than the original version, while performing weaker on ECA, EECS, and SECS. We guess that the synthesis quality and pronunciation are better due to the additional information encoded from the reference audio that helps with speech synthesis. However, the lack of style disentanglement leads to decreased performance on emotion and speaker-related metrics. The results of the unseen speaker show that the modified structure allows for the emotion conversion of a new speaker without losing quality. Fig. 10 shows the pitch tracks of the transformed samples of the unseen speaker for each emotion, showing distinct differences for each emotion. We found that the speaker similarity of unseen speakers was poor compared to seen speakers. Table XI shows the speaker similarity for each emotion. We observe a decrease in the speaker similarity for all emotions. Fig. 11 shows a t-SNE visualization of speaker embedding for the ESD dataset (green), the VCTK dataset (purple), and the results of the unseen speaker experiment (yellow). The model synthesized speech that was closer to the speaker in the ESD, which was the training set.
@@ V-C Unseen Speaker Emotion Conversion
VI Discussion
VI-A Limitations
In experimental results, diffusion-based model using demonstrated improved results. However, the limitations of diffusion-based structures are their extensive computational demand and time-consuming nature. We anticipate that this challenge will be alleviated by the advent of the recent fast sampling method[76]. Some applications [77] have been made in speech research. The scope of our experiments also encompassed speaker generalization. Although we successfully observed emotional transformations in the voices of unseen speakers, a discernible lack of speaker similarity was apparent, indicating the need for further refinement. Empirically, our observations suggest that models tested on unseen data often reflect the voice distribution of the data used for their pre-training. Expanding the dataset is expected to increase the representational capacity of the generator, and the performance of zero-shot emotion conversion is also expected to improve. In the configuration of our style autoencoder, we utilized MixLN for de-stylize and CLN for stylize. It was observed that the perturbations introduced by MixLN resulted in a compromise that affected the equilibrium between expressiveness and pronunciation accuracy within our model. In addition, the task of effectively separating style from content remains a significant challenge [78, 79], requiring ongoing research and development to address this issue. Although the diffusion model produces high-quality speech, it has limitations because emotion datasets typically have a 16k sampling rate. The audio super-resolution models such as [80] are expected to solve this problem. Although we only experimented with English, it has the potential to be extended by considering a wide range of languages and combining it with speech-to-speech translations [81]. The expression of emotions differs between people, languages, and cultures, and research is needed to reflect these differences.
@@ VI-B Future Works
Fundamentally, the proposed model depends on the performance of the SSL model because it utilizes discrete units. In addition to semantic units such as HuBERT, we also plan to investigate structures that exploit neural audio codecs such as [39, 38]. Recent research in the field of speech synthesis has increasingly focused on controlling emotion intensity. Some works, such as [82] and [83], have adopted the method of modeling emotion intensity using relative attribute ranking functions. Otherwise, some studies, such as [84], have explored the modeling of intensity through interpolation of embeddings. In addition, various approaches have tried to control the intensity of emotions, such as [85]. Although these studies have shown that modeling emotion intensity is possible, precise control of intensity remains a challenge. In our future work, we explore the direction in which emotional intensity can be controlled. We will also investigate emotional voice conversion for cross-language. Furthermore, it is expected to be applied to singing voice synthesis tasks [86] to express emotions.
@@ VII Conclusion
In this work, we introduced DurFlex-EVC, which generates speech of various durations. We leveraged the discrete speech units of HuBERT to model the contents at the unit-level, and achieved duration flexibility by predicting unit duration and extending it to the frame-level. We propose a style autoencoder to disentangle the source style of input features and apply target styles, and a unit aligner to enable emotional context modeling at the unit level. A hierarchical stylize encoder is introduced for stylistic enhancement of features. Further improvements were achieved by including a stochastic duration predictor and a diffusion-based generator. Experiments demonstrated that DurFlex-EVC outperforms existing EVC models in terms of performance. We also extended the models capabilities to scenarios with unseen speakers, and DurFlex-EVC effectively converted emotions while maintaining a high level of speech quality and pronunciation accuracy. We performed a variety of experiments to validate the effectiveness of our proposed model and believe that it can make a significant contribution to the advancement of emotional speech recognition.

### v4

@@ V-B3 Effectiveness of Hierarchical Stylize Encoder
V-B4 Comparison for Duration Predictors
We introduced a stochastic duration predictor (SDP) for duration modeling to represent the diversity of emotions. To evaluate, we compared our method with widely used deterministic duration predictors (DDP). The DDP follows the structure described in FastSpeech [24], which includes two convolutional layers, ReLU activation, layer normalization, and dropout. Tables V, w/ DDP shows the evaluation results of the model using DDP. The evaluation results show that the model using SDP is better than the model using DDP. We compared the JSD of unit duration for each emotion for each duration predictor. Table VII shows the JSD results for unit duration for models with stochastic duration predictors, w/ SDP, and deterministic duration predictors, w/ DDP.
(a) DurFlex-EVC w/ FFT
(b) DurFlex-EVC w/ diffusion
@@ V-B5 Comparison for Mel-spectrogram Generator Structures
To verify the expressiveness of each emotion, we computed the JSD over pitch for each emotion. In Table VIII, w/ FFT is the model using an FFT decoder and w/ diffusion is the model using a diffusion-based structure. w/ diffusion shows better JSD on neutral, angry, sad, while w/ FFT shows better JSD in happy and surprise. Although w/ FFT outscores w/ diffusion in Happy and Surprise, it does not mean that FFT is more expressive. Fig. 8 shows the pitch of the speech generated by multiple iterations of the same sentence. Models using FFT produce consistent pitch tracks that remain stable over repeated generations, while models using diffusion produce more dynamic results. This suggests that the diffusion-based structure captures higher expressiveness, resulting in a higher JSD compared to the FFT-based structure, which tends to average over dynamic emotions like happy and surprise.
Figure 9: Style embedding modeling for unseen speaker setting.
@@
V-B6 Comparison for Input Features
The proposed model takes HuBERT features as input and outputs a Mel-spectrogram. The reason for outputting Mel-spectrogram is for compatibility with pre-trained vocoders. To explain the reasoning behind our choice of HuBERT features, we present a comparison of the input features. We compared Mel-spectrogram and linear spectrogram as features using conventional signal processing, and wav2vec 2.0 [31], wavLM [70], and HuBERT [33] as SSL features that have recently been used as linguistic representations [71, 35].
Although the input feature dimensions differ in each experiment, they are aligned to the same hidden dimension through a linear layer.
Table IX shows the results of the experiments by input feature. The results show that models using Mel-spectrogram and linear-spectrogram have lower UTMOS than SSL feature models. Models using wavLM performed better overall than those using wav2vec 2.0. The model using HuBERT outperformed the others in all metrics except ECA and EECS. In particular, the model using HuBERT shows a significant improvement in pronunciation, which is interpreted as advantageous for linguistic learning over other features because the target units used for training are obtained from the clustering of HuBERT features. To compare the linguistic modeling ability of each input feature, we calculated the BLEU score [72] and the unit error rate (UER) of the predicted units. Table X shows the BLEU score and UER of the predicted units for each feature. We found that the prediction of units was correlated with the pronunciation accuracy of the synthesized speech, and this was due to the fact that HuBERT predicted units more accurately than the other features.
We also experimented with a model that takes unit input and generates speech. In Table V, w/ unit2mel is a model that receives unit input and generates Mel-spectrogram. This model has the same structure, with the addition of a unit embedding layer to accept units as input. Furthermore, we experimented with models that generate waveforms directly from unit. w/ unit2wav is a modification of HiFi-GAN [48] to take unit as input, which is equivalent to not including f0 in Textless-EVC generator.
The w/ unit2mel shows a worse overall performance, except for ECA and EECS. The w/ unit2wav shows significantly worse metrics across the board, except for CER and WER. We interpret this to mean that the speech unit has enough information about pronunciation but not enough other speech information to generate a waveform. To overcome this, Textless-EVC uses additional information such as pitch and timbre to generate the waveform.
@@ V-B6 Comparison for Input Features
V-C Unseen Speaker Emotion Conversion
Table XI shows the evaluation results for seen and unseen speakers for the modified model. The modified model scored better UTMOS, PER, CER, and WER than the original version, while performing weaker on ECA, EECS, and SECS. We infer that the synthesis quality and pronunciation are better due to the additional information encoded from the reference audio that helps with speech synthesis. However, the lack of style disentanglement leads to decreased performance on emotion and speaker-related metrics. The results of the unseen speaker show that the modified structure allows for the emotion conversion of a new speaker without losing quality. Fig. 10 shows the pitch tracks of the transformed samples of the unseen speaker for each emotion, showing distinct differences for each emotion. We found that the speaker similarity for unseen speakers was lower compared to seen speakers. Table XI shows the speaker similarity for each emotion. We observe a decrease in the speaker similarity for all emotions. Fig. 11 shows a t-SNE visualization of speaker embedding for the ESD dataset (green), the VCTK dataset (purple), and the results of the unseen speaker experiment (yellow). The model synthesized speech that was closer to the speaker in the ESD, which was the training set.
VI Discussion
VI-A Limitations
In experimental results, the diffusion-based model demonstrated improved results. However, the limitations of diffusion-based structures include their extensive computational demand and time-consuming nature. We anticipate that this challenge will be alleviated by the advent of the recent fast sampling method[76]. Some applications have been developed in speech research [77]. The scope of our experiments also encompassed speaker generalization. Although we successfully observed emotional transformations in the voices of unseen speakers, a discernible lack of speaker similarity was apparent, indicating the need for further refinement. Empirically, our observations suggest that models tested on unseen data often reflect the voice distribution of the data used for their pre-training. Expanding the dataset is expected to increase the representational capacity of the generator, and the performance of zero-shot emotion conversion is also expected to improve. In the configuration of our style autoencoder, we utilized MixLN for de-stylizing and CLN for stylizing. It was observed that the perturbations introduced by MixLN resulted in a compromise that affected the equilibrium between expressiveness and pronunciation accuracy within our model. In addition, the task of effectively separating style from content remains a significant challenge [78, 79], requiring ongoing research and development to address this issue.
Although the diffusion model produces high-quality speech, it has limitations due to emotion datasets typically having a 16k sampling rate. Audio super-resolution models, such as [80], are expected to solve this problem. Although we only experimented with English, our model has the potential to be extended by considering a wide range of languages and combining it with speech-to-speech translations [81]. The expression of emotions differs between people, languages, and cultures, and research is needed to reflect these differences.
@@ VI-B Future Works
Fundamentally, the proposed model depends on the performance of the SSL model because it utilizes discrete units. In addition to semantic units such as HuBERT, we also plan to investigate structures that exploit neural audio codecs such as [39, 38]. Recent research in the field of speech synthesis has increasingly focused on controlling emotion intensity. Some works, such as [82] and [83], have adopted the method of modeling emotion intensity using relative attribute ranking functions. Alternatively, some studies, such as [84], have explored the modeling of intensity through interpolation of embeddings. Various approaches, such as [85], have tried to control the intensity of emotions. Although these studies have shown that modeling emotion intensity is possible, precise control of intensity remains a challenge. In our future work, we will explore methods to control emotional intensity. We will also investigate emotional voice conversion for cross-language. Furthermore, it is expected to be applied to singing voice synthesis tasks [86] to express emotions.
@@ VII Conclusion
In this work, we proposed DurFlex-EVC, a robust emotional voice conversion framework that supports flexible durations without relying on text or alignment. Exploiting discrete speech units derived from HuBERT, our model encodes content at the unit level and allows flexible duration control by extending unit-level duration prediction to the frame level. We developed a style autoencoder to disentangle emotional style from content, allowing precise manipulation of target styles. Additionally, we introduced a unit aligner to model emotional context at the unit level. A hierarchical stylize encoder further enhances emotional expressiveness by refining style application at both unit and frame levels. To improve synthesis quality, we incorporated a stochastic duration predictor and a diffusion-based generator, ensuring naturalness and diversity in generated speech. Through extensive experiments, DurFlex-EVC demonstrated superior performance over existing models, excelling in both emotional reproduction accuracy and speech quality. The framework effectively generalized to unseen speakers, preserving pronunciation and emotional clarity. These results demonstrate its potential for advancing emotional voice conversion, particularly in applications requiring duration flexibility and expressive emotional transformations.